OpenStudy (anonymous):

WILL FAN AND MEDAL!!! can someone please help on like 5 questions of math please

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

http://prntscr.com/a5cdo3

OpenStudy (igreen):

You can use the substitution method.

OpenStudy (anonymous):

ok??? is that like put in a a number for x and y

OpenStudy (igreen):

Subtract 'y' to both sides in the first equation.

OpenStudy (anonymous):

ok

OpenStudy (igreen):

Tell me what you come up with.

OpenStudy (anonymous):

so x= 5-y

OpenStudy (anonymous):

or x =y-5

OpenStudy (igreen):

Yep! Now plug in '5 - y' for 'x' in the 2nd equation and solve for 'y'.

OpenStudy (igreen):

5 - y and y - 5 aren't the same.

OpenStudy (mathmale):

Your x=5-y is fine. Now, what are you going to do with this? Hint: substitution method.

OpenStudy (igreen):

Can you plug in '5 - y' for 'x' in the 2nd equation and solve for 'y'?

OpenStudy (anonymous):

im sorrry this makes o sense to me... lol

OpenStudy (igreen):

Okay, we have: x = 5 - y for the first equation Our second equation is: x - 2y = 2 Now in the first equation, x is equal to 5 - y, so we can plug that in for 'x' in the second equation. 5 - y - 2y = 2 Now solve that for 'y', can you do that?

OpenStudy (anonymous):

3y =-3 so y = -1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

???

OpenStudy (ineedhelplz):

x=4 and y=1

OpenStudy (xmissalycatx):

You have to combine like terms in 5 - y - 2y = 2. So it would actually be 5 - 3y = 2

OpenStudy (xmissalycatx):

And then subtract 5 from both sides and divide by -1.

OpenStudy (anonymous):

tahnks so much

OpenStudy (xmissalycatx):

@Ineedhelplz Please don't give direct answers.. :/

OpenStudy (anonymous):

can you help me on another one

OpenStudy (xmissalycatx):

I sure can :P

OpenStudy (anonymous):

great

OpenStudy (error1603):

Wooooo

OpenStudy (anonymous):

http://prntscr.com/a5cpn7

OpenStudy (xmissalycatx):

Oh poop I can't open prntscr links :(

OpenStudy (anonymous):

ok why not

OpenStudy (xmissalycatx):

My computer blocks access to that website.

OpenStudy (anonymous):

aww

OpenStudy (xmissalycatx):

:( darnit.

OpenStudy (anonymous):

aww

OpenStudy (error1603):

Are you at home?

OpenStudy (xmissalycatx):

No I am not. I use public computers. My home computers I use for other things. I do actually have to go in approximately 5 minutes so.

OpenStudy (anonymous):

you can help when you get home???

OpenStudy (xmissalycatx):

I don't get home for another few hours. I'm sure someone will be able to come by and help you sooner than that. :)

OpenStudy (anonymous):

ok